OpenStudy (anonymous):

Solve the system, using substitution. y = 1 - 5x and 5x + y = 8 no solution infinitely many solutions

OpenStudy (johnweldon1993):

See how that first equation tells you what y equals? Substitute that into your second equation... 5x + y = 8 you know y = 1 - 5x so 5x + (1 - 5x) = 8 5x + 1 - 5x = 8 5x - 5x + 1 = 8 1 = 8 is this a true statement? If so...it has infinitely many solutions if not...it has no solutions

OpenStudy (anonymous):

so im guessing no solution?

OpenStudy (johnweldon1993):

That would be correct